OpenStudy (anonymous):

what is 0.12.... as a fraction in simplest form

OpenStudy (anonymous):

easiest way to go about these types of problems is to make the easiest fraction you can first, then simplify. By multiplying by 100, you get that

OpenStudy (anonymous):

\[.12=\frac{ 12 }{ 100}\]

OpenStudy (anonymous):

\[\frac{ 12 }{ 100 }=\frac{ 6 }{ 50 }= \frac{3}{25}\]